Circular No. 6991

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dk IN UGC 139 J. Y. King, M. Modjaz, T. Shefler, E. Halderson, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dk was found and confirmed in unfiltered CCD images obtained on Aug. 19.4 (mag ∼ 17.6) and 20.4 UT (mag ∼ 17.0). The new object is located at α = 0^{h}14^{m}32^{s}.16, δ = −0^{o}44′10′′.9 (equinox 2000.0), which is ∼ 5′′.4 east and 3′′.1 north of the nucleus of UGC 139. A KAIT image obtained on Aug. 15.4 shows no star at this position (limiting mag ∼ 18.5).

COMET C/1998 P1 (WILLIAMS) Improved parabolic orbital elements from MPEC 1998-Q10:

T = 1998 Oct. 17.836 TT  ω = 294.^{o}466
Ω = 156.376 }2000.0
q = 1.14678 AU  i = 145.729
1998 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 15 14^{h}35.^{m}93 −59^{o}41.^{′}8 1.054 1.520 94.^{o}7 41.^{o}6 8.4
20 14 13.39 −54 22.7 1.141 1.472 86.0 43.3 8.5
25 13 59.65 −49 48.4 1.237 1.426 78.0 43.9 8.5
30 13 50.75 −45 56.9 1.336 1.382 70.6 43.6 8.5
Sept. 4 13 44.75 −42 42.0 1.436 1.341 63.7 42.4 8.6
9 13 40.55 −39 56.9 1.534 1.302 57.2 40.6 8.6
14 13 37.54 −37 35.5 1.628 1.267 51.1 38.2 8.6
19 13 35.33 −35 33.2 1.716 1.236 45.2 35.3 8.6
24 13 33.67 −33 45.7 1.797 1.209 39.7 32.0 8.6
29 13 32.38 −32 10.0 1.870 1.186 34.3 28.4 8.6
Oct. 4 13 31.35 −30 43.5 1.934 1.168 29.3 24.7 8.6

Visual m_{1} estimates: Aug. 12.42 UT, 8.4 (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 13.47, 8.6 (C. E. Drescher, Warrill View, Queensland, Australia, 0.25-m refl.); 15.82, 8.8 (J. G. S. Aguiar, Campinas, S.P, Brazil, 11×80 binoculars); 17.47, 8.3 (S. T. Rae, Palmerston North, New Zealand, 0.11-m refl.); 19.91, 8.5 (W. C. de Souza, São Paulo, Brazil, 11×80 binoc.); 20.23, 8.5 (S. J. O’Meara, Volcano, HI, 0.10-m refractor).

1998 August 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green